OpenStudy (anonymous):

use even and odd properties of trig functions to find the exact value of the expression, cos(-60) posted 30 mins ago,no help

OpenStudy (sidsiddhartha):

let f(x) be a function so f(x) will be even when\[\boxed {f(-x)=f(x)}\]and f(x) will be odd when\[\boxed{f(-x)=-f(x)}\]

OpenStudy (sidsiddhartha):

now cos(x) is a even function so by this property we can write\[\cos(-x)=\cos(x)\] ok? make sense

OpenStudy (sidsiddhartha):

finally we can write\[\cos(-60)=\cos(60)=\frac{ 1 }{ 2 }\] got it? @dabeezy

OpenStudy (anonymous):